OpenStudy (anonymous):

related rates help. Suppose x thousand watches can be sold at p dollars per watch where p+x^2 = 144. How fast is the demand changing when x=9, p=63 and the price per watch is increasing at the rate of $2/week? Please do a step by step explanation. The answer is dx/dt = -1/9 so lets see how they got that. thanks.

OpenStudy (anonymous):

I solved it .. in 10 seconds ;)

OpenStudy (anonymous):

\[ p+x^2 = 144 \] derivate this with respect to t \[ \frac{dp}{dt} +2x \frac{dx}{dt} = 0 \] substitute x=9 and dp/dt =2 you will have your answer.

OpenStudy (anonymous):

why can't i do this stuff? im so terrible. thanks

OpenStudy (anonymous):

Naah you are just fine ..